Ted left his home by car traveling at 40 mph. His father left two hours later traveling in the same direction at 55 mph. How many hour will it take his father to overtake him?
Answer by ankor@dixie-net.com(22740)

Ted left his home by car traveling at 40 mph. His father left two hours later traveling in the same direction at 55 mph. How many hours will it take is father to overtake him?:
:
When Dad overtakes Ted, they will have traveled the same distance so it is easy to make a distance equation: Dist = speed*time
:
Let t = Dad's travel time when they meet: then (t+2) = Ted's travel time
:
Dad dist = Ted's Dist
55t = 40(t+2)
55t = 40t + 80
55t = 40t = 80
15t = 80
t = 80/15
t = 5 & 1/3 hrs: 5 hr 20 min
:
Check solution by finding the distance:
55*5.33 = 293.2 mi
40*7.33 = 293.2 mi also